# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ 4 Interpretation as a model for dark matter
Cosmological first-order phase transitions and the related bubble-nucleation processes have been studied extensively since the 1980’s [235, 236, 208, 237, 238, 239, 207, 240, 241, 242]. To obtain an approximate quantitative description of bubble nucleation at the phase transition, one can follow the same arguments as in ref. [67] (that we reproduce here, for the readers’ convenience), applying them to the $\operatorname{\mathrm{Sp}}(2)$ theory. For the $\operatorname{\mathrm{SU}}(3)$ theory, the latent heat and the interface tension $\sigma_{\mbox{\tiny{cd}}}$ between the deconfined phase and the confining phase were estimated in ref. [188], where it was found that $\frac{L_{\mbox{\tiny{h}}}}{\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}}}\simeq 70$; one can assume that this ratio takes a similar value in the $\operatorname{\mathrm{Sp}}(2)$ theory, too. This order-of-magnitude assumption is corroborated by the observation that, combining our result for the latent heat with the result for the $\frac{\sigma_{\mbox{\tiny{cd}}}}{T_{\mbox{\tiny{c}}}^{3}}$ ratio in the $\operatorname{\mathrm{Sp}}(2)$ theory, as obtained from simulations at a finite lattice spacing corresponding to $N_{\tau}=4$, that was recently reported in ref. [139], namely $\left.\frac{\sigma_{\mbox{\tiny{cd}}}}{T_{\mbox{\tiny{c}}}^{3}}\right|_{N_{\tau}=4}=0.01718(32)$, one obtains a roughly comparable estimate for the $\frac{L_{\mbox{\tiny{h}}}}{\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}}}$ ratio. Even though the result for the $\frac{\sigma_{\mbox{\tiny{cd}}}}{T_{\mbox{\tiny{c}}}^{3}}$ reported in ref. [139] is not extrapolated to the continuum limit, and corrections due to the finiteness of the lattice spacing in simulations with $N_{\tau}=4$ may be non-negligible, we take this as an indication of the validity of our assumption about the value of the $\frac{L_{\mbox{\tiny{h}}}}{\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}}}$ ratio in the $\operatorname{\mathrm{Sp}}(2)$ theory.
@@
Assume that, while the temperature decreases, the deconfined phase supercools down to a dark-sector temperature $T=(1-\delta)T_{\mbox{\tiny{c}}}$, with $\delta$ a small, positive real number. [^7] A small value of $\delta$ was indeed suggested, for example, by early studies modelling the cosmological change of phase in the strong-interaction sector of the Standard Model as a first-order transition [243, 244]. Then, if the confining phase sets in within a small region of space (a bubble), this region can start growing, as long as the force due to the difference in pressure between the two phases overcomes the force due to the interface tension between the two phases. For a spherically symmetric bubble, this occurs when its radius $r$ exceeds a critical value $r_{\mbox{\tiny{c}}}=\frac{2\sigma_{\mbox{\tiny{cd}}}}{\Delta p}$, where $\Delta p$ denotes the difference between the pressure $p_{\mbox{\tiny{c}}}(T)$ of the confining phase at temperature $T$ and the pressure of the deconfined phase $p_{\mbox{\tiny{d}}}(T)$, at the same temperature. Note that, since $\delta$ is small, $\Delta p$ can be approximated by working out the Taylor expansions of $p_{\mbox{\tiny{c}}}(T)$ and $p_{\mbox{\tiny{d}}}(T)$ around $T_{\mbox{\tiny{c}}}$ at the leading non-trivial order:
@@
where we denoted the $\frac{L_{\mbox{\tiny{h}}}}{\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}}}$ ratio for the $\operatorname{\mathrm{Sp}}(2)$ theory as $\xi$. Note that in the denominator of the last fraction in eq. (4.11) there is a potential competition (or at least partial compensation) between the third power of the large term $\xi$ and the square of the small parameter $\delta$. Using our estimate for $\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}$ given in eq. (3.13), and working under the aforementioned assumption $\xi\simeq 70$, eq. (4.11) yields
@@
Recalling that the bubble nucleation rate per unit volume can be related to the free-energy cost of a bubble of critical radius via [246] (for a detailed discussion, see also ref. [196, section 9.2])
$\frac{\Gamma T_{\mbox{\tiny{c}}}^{4}}{V}\simeq\exp\left(-\frac{\Delta F}{T_{\mbox{\tiny{c}}}}\right),$ (4.13)
one finds that $\frac{\Gamma}{V}\simeq H^{4}$ (corresponding, on average, to one bubble nucleation per Hubble time taking place within one Hubble volume) when
@@
It is interesting to consider the implications of eq. (4.14) for the (unphysical) case of first-order transitions in the Standard Model, through a back-of-the-envelope calculation based on the values for the Hubble radius reported in ref. [245]. If the electroweak crossover were a first-order transition, then a crude estimate for the logarithm on the right-hand side of eq. (4.14) would be of the order of $37$; similarly, if the QCD crossover were a first-order transition, then the logarithm on the right-hand side of eq. (4.14) would be of the order of $44$. Thus, in both cases the $\frac{\Delta F}{T_{\mbox{\tiny{c}}}}$ ratio would be rather large, corresponding to small values of $\delta$, of the order of $O(10^{-4})$, implying that even for modest supercooling the heat released by the formation and expansion of bubbles would increase again the temperature above the critical one, except if the transition can be completed within a very small fraction of a Hubble time [^8] For a first-order transition in the strong-interaction sector of the Standard Model, this could happen if the heat capacity of the quark-gluon plasma was sufficiently large, which would allow the supercooled deconfined medium to absorb the heat released by a bubble in its expansion. [67].
@@
This reasoning, however, cannot be directly extended to our dark $\operatorname{\mathrm{Sp}}(2)$ glueball model, given that a priori there is no obvious reason to assume that its temperature coincides with the one of the visible sector, and thus, in particular, there is no compelling argument allowing one to derive an estimate for the right-hand side of eq. (4.14). As a consequence, if supercooling down to temperatures not too close to $T_{\mbox{\tiny{c}}}$ could take place in our model, then a non-negligible amount of entropy could be produced during the bubble-growth process, in which the system is out of thermodynamic equilibrium. This could potentially lead to a detectable effect on relic densities. Note that in this respect our model is different from the “twin $\operatorname{\mathrm{SU}}(3)$” model that was discussed in ref. [67].
@@
After the transition to the confining phase is complete and the temperature continues to decrease, the dark sector will behave as a gas of massive, weakly interacting glueballs. Note that the kinetic-equilibrium condition implies that, for a generic glueball of mass $m$ at temperature $T$ and for chemical potential $\mu$, the particle density is expected to be
@@
Note that in the non-relativistic limit eq. (4.15) reduces to
@@
Depending on phase space and kinematics, at this stage heavier dark glueballs may decay to lighter ones, until a freeze-out condition is reached. In addition, even in the absence of direct couplings to SM fields (a particularly intriguing scenario, in which the characteristic mass scale of dark matter could evade current lower bounds from collider experiments [104, 247]) the dark sector can be connected to the visible sector via non-renormalizable operators [63, 64], which can transfer energy and possibly allow dark glueballs (at least those in some channels) to decay. The existence of such “connector operators” is generic, as some degree of mixing between the visible and dark sectors is expected to be induced at least by quantum gravity effects, but in this case the mixing between sectors would be suppressed by some power of the ratio between the dark confinement scale and the Planck scale [248]. In particular, dark glueballs are expected to decay to gravitons, with a decay to two gravitons being the dominant channel, but the corresponding decay rate would scale proportionally to the inverse of the fourth power of the Planck scale and like the fifth power of the glueball mass. This would imply a glueball lifetime exceeding the age of the universe for glueball masses below $10^{7}~{}\mathrm{GeV}$ [62]. By contrast, a stronger mixing could appear in the presence of additional fields, directly coupled to both sectors. Depending on the strength of the couplings of the connector operators, dark glueballs decaying to the visible sector could then yield potentially observable contributions to the electromagnetic radiation emitted by astrophysical sources [249], to the cosmic microwave background [250], and to Big-Bang nucleosynthesis [251, 252, 253].
If the mixing between the visible and dark sectors due to connector operators is negligible, then the two sectors can evolve with different temperatures. In particular, the relic dark matter abundance may have been determined by asymmetric reheating, whereby the same mechanism that determines the SM abundances in the early universe is at play for the dark sector, too, albeit at a different temperature [254]. Generalizing the discussion for a dark $\operatorname{\mathrm{SU}}(N)$ Yang–Mills theory presented in ref. [72] to the $\operatorname{\mathrm{Sp}}(2)$ model, the ratio $R$ of the entropy density of the dark sector over the entropy density of the SM (whose value is determined by the details of the reheating process) is then expected to remain constant during the evolution immediately after reheating; its value is related to the temperatures of the two sectors via
@@
Equation (4.17) is expected to hold at temperatures $T_{\mbox{\tiny{dark}}}$, that are sufficiently high, and well above the temperature of the confinement/deconfinement transition of the dark sector. When the temperature of the dark sector is reduced to $T_{\mbox{\tiny{c}}}$, the entropy density has a discontinuity, $s|_{T=T_{\mbox{\tiny{c}}}^{-}}=s|_{T=T_{\mbox{\tiny{c}}}^{+}}-\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}}$, and the system enters the confining phase, in which its equilibrium-thermodynamics can be described in terms of a gas of glueballs, according to eq. (3.8).
As the system evolves further, the temperature continues to decrease, with dark glueball interactions taking place; these interactions, in particular, include the annihilation of heavier glueballs, which occurs primarily through $2\to 2$ decays [68, 255], as well as the self-annihilation of the lightest glueball, e.g., via $3\to 2$ processes. Such interactions continue until freeze-out, which takes place when the rate of $3\to 2$ self-annihilation processes becomes too small compared to the Hubble expansion. From that point, the densities of the various dark glueball species remain fixed and simply dilute with the expansion of the universe, and $2\to 2$ elastic scatterings (which, from general arguments based on effective field theory, are expected to be parametrically favored with respect to $3\to 2$ decays) maintain kinetic equilibrium, so that stable relic densities are obtained.
For a dark $\operatorname{\mathrm{SU}}(N)$ Yang–Mills theory (with $N>2$), in ref. [70] it was shown that the time evolution of the densities of the lightest glueball species in the spectrum, corresponding to lowest-mass states in the $J^{PC}=0^{++}$ and $J^{PC}=1^{+-}$ channels, can be approximated well by two coupled equations, relating them to each other via the thermally averaged cross sections for glueball interactions. For our $\operatorname{\mathrm{Sp}}(2)$ model, however, the discussion has to be modified, as the theory features no $C=-1$ states. Restricting, for simplicity, the analysis to the lightest glueball state only, the associated particle density $n_{0^{+}}$ is expected to evolve in time according to
@@
In passing, it is worth noting that ref. [256] recently discussed the spectrum of gravitational waves that one can obtain in a theory with a Hagedorn-like spectrum. This may be of relevance for our model, too, since our results for the equation of state in the confining phase do provide evidence for an exponential growth of the density of states at temperatures close to $T_{\mbox{\tiny{c}}}$.
@@
Finally, we point out some comments on the possible values of the temperature of the dark sector, which is directly relevant for the existence of a phase transition, for the details of the spectrum of gravitational waves that the latter can produce, as well as for the freeze-out yields of candidate dark-matter states. The issue was recently discussed in ref. [87], which pointed out various important aspects. Firstly, the very definition of a “temperature” for the dark sector implicitly assumes that thermodynamic equilibrium can be reached; for a non-Abelian gauge theory, the equilibration rate is expected to be parametrically of the form $g^{2}T$, and this quantity should be larger than the Hubble rate. In turn, the latter is determined by the square of the largest, between the temperature of the visible sector and the temperature of the dark sector, divided by the Planck mass, so that the dark sector can reach thermal equilibrium for temperatures up to $g^{2}m_{\mbox{\tiny{Planck}}}$, with $m_{\mbox{\tiny{Planck}}}$ being the Planck mass. Secondly, it was shown that more constraining bounds on the allowed temperatures for the dark sector can be obtained, by assuming a “natural inflation” scenario [257], whereby an axion-like inflaton (with a simple cosine potential with parameters determined from experimental constraints [258]) directly couples to the dark sector: the reasoning starts from the observation that the smallness of equilibrium thermodynamic functions in the confining phase implies that even a modest energy transfer from the inflaton field can lead to a significant increase of the dark temperature. By solving the equations determining the time evolution of the plasma temperature and of the average value of the inflaton, the authors of ref. [87] found that if the ratio of the confining scale of dark sector over $m_{\mbox{\tiny{Planck}}}$ is $O(10^{-8})$ or larger, then the deconfinement temperature may not be accessible to the dark sector. Thirdly, they found that, for sufficiently high temperatures, the thermal, monotonically increasing intermediate-frequency contribution to the gravitational-wave background discussed in ref. [259] may be close to the detectability range of forthcoming experiments. Finally, they also discussed the gravitational waves produced at the confinement/deconfinement phase transition, finding that the peak frequencies predicted by their calculations could be observable by the Einstein telescope and/or DECIGO. [^9] A further, detailed discussion about this subject was later presented in ref. [260]. While the discussion of ref. [87] was carried out for a dark $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory, in view of the many similarities with the $\operatorname{\mathrm{Sp}}(2)$ theory, one may argue that their results can be expected to be qualitatively (and probably also semi-quantitatively) relevant for our model, too.

### v2

@@ 4 Interpretation as a model for dark matter and gravitational-wave production
Cosmological first-order phase transitions and the related bubble-nucleation processes have been studied extensively since the 1980’s [325, 326, 249, 327, 328, 329, 248, 330, 331, 332]. To obtain an approximate quantitative description of bubble nucleation at the phase transition, one can follow the same arguments as in ref. [77] (that we reproduce here, for the readers’ convenience), applying them to the $\operatorname{\mathrm{Sp}}(2)$ theory. For the $\operatorname{\mathrm{SU}}(3)$ theory, the latent heat and the interface tension $\sigma_{\mbox{\tiny{cd}}}$ between the deconfined phase and the confining phase were estimated in ref. [216], where it was found that $L_{\mbox{\tiny{h}}}/(\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}})\simeq 70$; one can assume that this ratio takes a similar value in the $\operatorname{\mathrm{Sp}}(2)$ theory, too. This order-of-magnitude assumption is corroborated by the observation that, combining our result for the latent heat with the result for the $\sigma_{\mbox{\tiny{cd}}}/T_{\mbox{\tiny{c}}}^{3}$ ratio in the $\operatorname{\mathrm{Sp}}(2)$ theory, as obtained from simulations at a finite lattice spacing corresponding to $N_{\tau}=4$, that was reported in ref. [153], namely $(\sigma_{\mbox{\tiny{cd}}}/T_{\mbox{\tiny{c}}}^{3})_{N_{\tau}=4}=0.01718(32)$, one obtains a roughly comparable estimate for the $L_{\mbox{\tiny{h}}}/(\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}})$ ratio. Even though the result for the $\sigma_{\mbox{\tiny{cd}}}/T_{\mbox{\tiny{c}}}^{3}$ reported in ref. [153] is not extrapolated to the continuum limit, and corrections due to the finiteness of the lattice spacing in simulations with $N_{\tau}=4$ may be non-negligible, we take this as an indication of the validity of our assumption about the value of the $L_{\mbox{\tiny{h}}}/(\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}})$ ratio in the $\operatorname{\mathrm{Sp}}(2)$ theory.
@@
Assume that, while the temperature decreases, the deconfined phase supercools down to a dark-sector temperature $T=(1-\delta)T_{\mbox{\tiny{c}}}$, with $\delta$ a small, positive real number. [^9] A small value of $\delta$ was indeed suggested, for example, by early studies modelling the cosmological change of phase in the strong-interaction sector of the Standard Model as a first-order transition [333, 334]. Then, if the confining phase sets in within a small region of space (a bubble), this region can start growing, as long as the force due to the difference in pressure overcomes the force due to the interface tension between the two phases. For a spherically symmetric bubble, this occurs when its radius $r$ exceeds a critical value $r_{\mbox{\tiny{c}}}=2\sigma_{\mbox{\tiny{cd}}}/(\Delta p)$, where $\Delta p$ denotes the difference between the pressure $p_{\mbox{\tiny{c}}}(T)$ of the confining phase at temperature $T$ and the pressure of the deconfined phase $p_{\mbox{\tiny{d}}}(T)$, at the same temperature. Note that, since $\delta$ is small, $\Delta p$ can be approximated by working out the Taylor expansions of $p_{\mbox{\tiny{c}}}(T)$ and $p_{\mbox{\tiny{d}}}(T)$ around $T_{\mbox{\tiny{c}}}$ at the leading non-trivial order:
@@
where we denoted the $L_{\mbox{\tiny{h}}}/(\sigma_{\mbox{\tiny{cd}}}T_{\mbox{\tiny{c}}})$ ratio for the $\operatorname{\mathrm{Sp}}(2)$ theory as $\xi$. Note that in the denominator of the last fraction in eq. (4.18) there is a potential competition (or at least partial compensation) between the third power of the large term $\xi$ and the square of the small parameter $\delta$. Using our estimate for $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}$ given in eq. (3.13), and working under the aforementioned assumption $\xi\simeq 70$, eq. (4.18) yields
@@
Recalling that the bubble nucleation rate per unit time and unit volume can be related to the free-energy cost of a bubble of critical radius via [336]
$\Gamma\simeq T_{\mbox{\tiny{c}}}^{4}\exp\left(-\frac{\Delta F}{T_{\mbox{\tiny{c}}}}\right)$ (4.20)
(for a detailed discussion, see also ref. [226, section 9.2]), one finds that $\Gamma\simeq H^{4}$ (corresponding, on average, to one bubble nucleation per Hubble time taking place within one Hubble volume) when
@@
It is interesting to consider the implications of eq. (4.21) for the (unphysical) case of first-order transitions in the Standard Model, through a back-of-the-envelope calculation based on the values for the Hubble radius reported in ref. [335]. If the electroweak crossover were a first-order transition, then a crude estimate for the logarithm on the right-hand side of eq. (4.21) would be of the order of $37$; similarly, if the QCD crossover were a first-order transition, then the logarithm on the right-hand side of eq. (4.21) would be of the order of $44$. Thus, in both cases the $\frac{\Delta F}{T_{\mbox{\tiny{c}}}}$ ratio would be rather large, corresponding to small values of $\delta$, of the order of $O(10^{-4})$, implying that even for modest supercooling the heat released by the formation and expansion of bubbles would increase again the temperature above the critical one, unless the transition could be completed within a very small fraction of a Hubble time [^10] For a first-order transition in the strong-interaction sector of the Standard Model, this could happen if the heat capacity of the quark-gluon plasma was sufficiently large, which would allow the supercooled deconfined medium to absorb the heat released by a bubble in its expansion. [77].
@@
This reasoning, however, cannot be directly extended to our dark $\operatorname{\mathrm{Sp}}(2)$ glueball model, given that a priori there is no obvious reason to assume that its temperature coincides with the one of the visible sector, and thus, in particular, there is no compelling argument allowing one to derive an estimate for the right-hand side of eq. (4.21). As a consequence, if supercooling down to temperatures not too close to $T_{\mbox{\tiny{c}}}$ could take place in our model, then a non-negligible amount of entropy could be produced during the bubble-growth process, in which the system is out of thermodynamic equilibrium. This could potentially lead to a detectable effect on relic densities. [^11] The non-equilibrium nature of bubble nucleation and expansion can be related with a wide variety of interesting phenomenological implications [337]. For example, in refs. [309, 338, 339] it was pointed out that, when the bubble wall velocity is sufficiently large, a non-thermal dark-matter production mechanism can set in, which could lead to a large relic abundance of particles with masses much larger than the scale of the transition, or to a non-trivial baryon-asymmetry generation. Note that in this respect our model is different from the “twin $\operatorname{\mathrm{SU}}(3)$” model that was discussed in ref. [77].
Glueballs in the confining phase as dark matter. At temperatures lower than $T_{\mbox{\tiny{c}}}$, the thermodynamics of the $\operatorname{\mathrm{Sp}}(2)$ gauge theory can be modelled in terms of a gas of massive glueballs: after confinement, a freeze-out process takes place, leading to a gas dominated by the contribution of the lightest state in the spectrum, i.e., the scalar glueball. If the dynamics of the latter can be described in terms of an effective Lagrangian of the form
which accounts for number-changing glueball self-interaction processes, like $3\to 2$ scatterings, then the number density of the glueball is expected to decrease as a function of time. While the $\lambda_{n}$ coefficients appearing in eq. (4.22) are a priori unknown, one can extract information about them by simple arguments. In particular, from naïve dimensional analysis follows that the energy dimension of the generic $\lambda_{n}$ coefficient is 4n4-n. A slightly less trivial observation is that, since $\phi$ represents a glueball state in a Yang-Mills theory in which the “color” multiplicity of the dark gluons is $10$, one could invoke large-$N$ arguments [340] (see also ref. [341, section 2.2]) to argue that $\lambda_{n}=O(N^{2-n})$, so that the dominant glueball self-interaction processes are encoded in the cubic term appearing in eq. (4.22). As long as the self-interaction rate is larger than the expansion rate of the universe, the comoving entropy density of the gas of dark glueballs is conserved. Noting that the results for the glueball masses in table 1 and the value of the $\frac{T_{\mbox{\tiny{c}}}}{\sqrt{\sigma}}$ ratio reported in ref. [35] indicate that in the confining phase one can use a non-relativistic approximation for the equation of state, and including only the leading contribution for the lightest glueball (denoting its mass as $m$), the entropy density reduces to
$s\simeq m^{2}\sqrt{\frac{mT}{8\pi^{3}}}\exp\left(-\frac{m}{T}\right).$ (4.23)
As a consequence, entropy conservation in a comoving volume $a^{3}$ implies that the temperature of dark matter decreases only logarithmically with $a$:
$T\simeq\frac{m}{3\ln(a/a_{0})},$ (4.24)
@@
Kinetic equilibrium implies that, for a generic glueball of mass $m$ at temperature $T$ and for chemical potential $\mu$, the particle density is expected to be
@@
In the non-relativistic limit eq. (4.25) reduces to
@@
Depending on phase space and kinematics, at this stage heavier dark glueballs may decay to lighter ones, until a freeze-out condition is reached. In addition, even in the absence of direct couplings to SM fields (a particularly intriguing scenario, in which the characteristic mass scale of dark matter could evade current lower bounds from collider experiments [118, 342]), the dark sector can be connected to the visible sector via non-renormalizable operators [73, 74], which can transfer energy and possibly allow dark glueballs (at least those in some channels) to decay. The existence of such “connector operators” is generic, as some degree of mixing between the visible and dark sectors is expected to be induced at least by quantum-gravity effects, but in this case the mixing between sectors would be suppressed by some power of the ratio of the dark confinement scale to the Planck scale [343]. In particular, dark glueballs are expected to decay to gravitons, with a decay to two gravitons being the dominant channel, but the corresponding decay rate would scale proportionally to the inverse of the fourth power of the Planck scale and like the fifth power of the glueball mass. Following the arguments of ref. [72], this would imply a glueball lifetime exceeding the age of the universe for glueball masses below $10^{7}~\mathrm{GeV}$. By contrast, a stronger mixing could appear in the presence of additional fields, directly coupled to both sectors. Depending on the strength of the couplings of the connector operators, dark glueballs decaying to the visible sector could then yield potentially observable contributions to the electromagnetic radiation emitted by astrophysical sources [344], to the cosmic microwave background [345], and to Big-Bang nucleosynthesis [346, 347, 348].
If the mixing between the visible and dark sectors due to connector operators is negligible, then the two sectors can evolve with different temperatures. In particular, the relic dark matter abundance may have been determined by asymmetric reheating, whereby the same mechanism that sets the SM abundances in the early universe is at play for the dark sector, too, but at a different temperature [349]. Generalizing the discussion for a dark $\operatorname{\mathrm{SU}}(N)$ Yang–Mills theory presented in ref. [83] to the $\operatorname{\mathrm{Sp}}(2)$ model, the ratio $R$ of the entropy density of the dark sector to the entropy density of the SM (whose value is determined by the details of the reheating process) is then expected to remain constant during the evolution immediately after reheating; its value is related to the temperatures of the two sectors via
@@
Equation (4.27) is expected to hold at temperatures $T_{\mbox{\tiny{dark}}}$, that are sufficiently high, and well above the temperature of the confinement/deconfinement transition of the dark sector. However, as we discussed, when the temperature of the dark sector is reduced to $T_{\mbox{\tiny{c}}}$, the entropy density has a discontinuity, $s|_{T=T_{\mbox{\tiny{c}}}^{-}}=s|_{T=T_{\mbox{\tiny{c}}}^{+}}-L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}$, and the system enters the confining phase, in which its equilibrium-thermodynamics can be described in terms of a gas of glueballs, according to eq. (3.8).
As the system evolves further, the temperature continues to decrease, with dark glueball interactions taking place; these interactions, in particular, include the annihilation of heavier glueballs, which occurs primarily through $2\to 2$ decays [78, 350], as well as the self-annihilation of the lightest glueball, e.g., via $3\to 2$ processes. Such interactions continue until freeze-out, which takes place when the rate of $3\to 2$ self-annihilation processes becomes too small compared to the Hubble expansion rate. From that point, the densities of the various dark glueball species remain fixed and simply dilute with the expansion of the universe, and $2\to 2$ elastic scatterings (which, from general arguments based on effective field theory, are expected to be parametrically favored with respect to $3\to 2$ decays) maintain kinetic equilibrium, so that stable relic densities are obtained.
For a dark $\operatorname{\mathrm{SU}}(N)$ Yang–Mills theory (with $N>2$), in ref. [80] it was shown that the time evolution of the densities of the lightest glueball species in the spectrum, corresponding to lowest-mass states in the $J^{PC}=0^{++}$ and $J^{PC}=1^{+-}$ channels, can be approximated well by two coupled equations, relating them to each other via the thermally averaged cross sections for glueball interactions. For our $\operatorname{\mathrm{Sp}}(2)$ model, however, the discussion has to be modified, as the theory features no C=1C=-1 states. Restricting, for simplicity, the analysis to the lightest glueball state only, the associated particle density $n_{0^{+}}$ is expected to evolve in time according to
@@
In passing, it is worth noting that refs. [351, 352] discussed the spectrum of gravitational waves that one can obtain in a theory with a Hagedorn-like spectrum. This may be of relevance for our model, too, since our results for the equation of state in the confining phase do provide evidence for an exponential growth of the density of states at temperatures close to $T_{\mbox{\tiny{c}}}$.
@@
Comments on the parameter space of the model. Finally, we point out some comments on the parameter space of the theory, including, in particular, on the possible values of the temperature of the dark sector, which is directly relevant for the existence of a phase transition, for the details of the spectrum of gravitational waves that the latter can produce, as well as for the freeze-out yields of candidate dark-matter states. The issue was discussed in ref. [100], which pointed out various important aspects. Firstly, the very definition of a “temperature” for the dark sector implicitly assumes that thermodynamic equilibrium can be reached; for a non-Abelian gauge theory, the equilibration rate is expected to be parametrically of the form $g^{2}T$, and this quantity should be larger than the Hubble rate. In turn, the latter is determined by the square of the largest, between the temperature of the visible sector and the temperature of the dark sector, divided by the Planck mass, so that the dark sector can reach thermal equilibrium for temperatures up to $g^{2}m_{\mbox{\tiny{Planck}}}$, with $m_{\mbox{\tiny{Planck}}}$ being the Planck mass. Secondly, it was shown that more constraining bounds on the allowed temperatures for the dark sector can be obtained, by assuming a “natural inflation” scenario [353], whereby an axion-like inflaton (with a simple cosine potential with parameters determined from experimental constraints [354]) directly couples to the dark sector: the reasoning starts from the observation that the smallness of equilibrium thermodynamic functions in the confining phase implies that even a modest energy transfer from the inflaton field can lead to a significant increase of the dark-sector temperature. By solving the equations determining the time evolution of the plasma temperature and of the average value of the inflaton, the authors of ref. [100] found that if the ratio of the confining scale of the dark sector to $m_{\mbox{\tiny{Planck}}}$ is $O(10^{-8})$ or larger, then the deconfinement temperature may not be accessible to the dark sector. Thirdly, they found that, for sufficiently high temperatures, the thermal, monotonically increasing intermediate-frequency contribution to the gravitational-wave background discussed in ref. [355] may be close to the detectability range of forthcoming experiments. Finally, they also discussed the gravitational waves produced at the confinement/deconfinement phase transition, finding that the peak frequencies predicted by their calculations could be observable by the Einstein Telescope and/or DECIGO. [^12] A further, detailed discussion about this subject was later presented in ref. [356]. While the discussion of ref. [100] was carried out for a dark $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory, in view of the many similarities with the $\operatorname{\mathrm{Sp}}(2)$ theory, one may argue that their results can be expected to be qualitatively (and probably also semi-quantitatively) relevant for our model, too.